The lesser of two numbers is 2 less than the greater number. If 10 is subtracted from six times the lesser number, the result is the same as 6 less than four times the greater number. Find the numbers.
Answer by mananth(16946)

x lesser number
bigger number =y
x=y-2
6x-10=4y-6
6x-4y=4
substitue x
6(y-2)-4y=4
6y-12-4y=4
2y=16
y=8
x=6
m.ananth@hotmail.ca